what is the answer to 11x-7x-24=32..... i was trying and i got many different answers
Answer by Maths68(1474)

11x-7x-24=32
4x-24=32
4x=32+24
4x=56
x=56/4
x=14
Check
Put value of x in the original equation
11x-7x-24=32
11(14)-7(14)-24=32
154-98-24=32
154-122=32
32=32